OpenStudy (anonymous):

A circular piece of metal has a radius of 3.5 meters. The metal sells for $3.25 per square meter. What is the total cost of the circular piece of metal? A. $11.38 B. $38.47 C. $71.44 D. $125.01

OpenStudy (anonymous):

I need Help.

OpenStudy (anonymous):

T o find how many meters square a circle has, you need to first find the formula for area of a circle area of a circle = Pie*r^2 so you input your data as follows A = 3.14 * (3.5 m)^2 A = 38.465 m^2 <--(meters squared) then you need to find how much it would be for each of these squared meters $$$ = 38.465 * $3.25 $$$ = 125.01125 = $125.01 so I would pick D

OpenStudy (anonymous):

Ok :D

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

if you need help on anymore just tag me